OpenStudy (anonymous):

If (-3,4) is a point on the graph of y=f(x), find b such that (5,b) is on the graph of y=f(x-8)

OpenStudy (anonymous):

oh its pretty simple. What happens if you plug 5 in for x in y=f(x-8)?

OpenStudy (anonymous):

You get y=f(-3)

OpenStudy (anonymous):

Right, so is it obvious to you that from the point (-3,4) on the graph of y=f(x) they plugged in -3 into x in the equation y=f(x)?

OpenStudy (anonymous):

I don't understand what you mean.

OpenStudy (anonymous):

I am talking about the first part of the question, when they plugged -3 in for x for the equation y=f(x), they got a y value of 4. and then if you plug in 5 into x in the equation y=f(x-8). you get y=f(-3)? see similarity there?

OpenStudy (anonymous):

and they want b, which is just the y value.

OpenStudy (anonymous):

do you understand?

OpenStudy (anonymous):

I don't understand how they got a value of 4 though and so the answer would be -3? And no I'm still confused.